Hayhayz (hayhayz):

Larry and Peggy are making decisions about their bank accounts. Larry wants to deposit $350 as a principle amount, with an interest of 4% compounded quarterly. Peggy wants to deposit $350 as the principle amount, with an interest of 6% compounded monthly. Explain which method results in more money after 2 years. Show all work. @imqwerty

imqwerty (imqwerty):

we use this formula to find the amount of money that we will get at the end- \(A=P(1+\large \frac{r}{n} )^{nt}\) okay here A is the amount that we will be getting P=the principle amount tat we deposited at the beginning n=number of months for which the principle is compounded r=rate of interest in decimals t=time for which the money is deposited so like lets find the amount that Larry will be getting- P=350 t=2years r=4%=0.04 it is quarterly compounded so 4months therefore n=4 put all this in the equation we get this- \(A=350(1+\large \frac{0.04}{4})^{4 \times 2}\)

Hayhayz (hayhayz):

Okay http://prntscr.com/af7tky

imqwerty (imqwerty):

yes now you need to find the amount ffor peggy

Hayhayz (hayhayz):

what would n be

Hayhayz (hayhayz):

p=350 t=2 years r=0.06

imqwerty (imqwerty):

yes and what bout n and t?

Hayhayz (hayhayz):

t is 2 years idk what n is

imqwerty (imqwerty):

well since the amount is compounded monthly n is 12

Hayhayz (hayhayz):

a=395

imqwerty (imqwerty):

yes correct

imqwerty (imqwerty):

now you can compare them to get your answer

Hayhayz (hayhayz):

so peggy method results in more money .. 2 more questions!

imqwerty (imqwerty):

yes :) yay :D